OpenStudy (anonymous):

Evaluate the limit and justify each step by indicating the appropriate limit laws lim (3x^4 + 2x^2 - x + 1 ) as x approaches -2

OpenStudy (ybarrap):

lim(3x^4 + 2x^2 - x + 1 ) = lim 3x^4 + lim 2x^2 - lim x + lim 1, lim of sum is sum of lim = lim3 * (lim x)^4 + lim2*(limx)^2 + lim(-1) limx + lim1, lim of product is product of lim, lim of exponent is exponent of lim = 3 (-2)^4 + 2 *(-2)^2 - (-2) + 1, evaluate as x -> -2, lim of constant is constant = 3*16 + 2*4 + 3 = 48 + 11 = 59

OpenStudy (anonymous):

i have some otehr limit problems..can u help me w/ those

OpenStudy (ybarrap):

post them to all, others may get to them sooner

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

i posted it